OpenStudy (anonymous):

in the function y+3=(1/3)^2+1 what does the 1/3 have on the graph compared toy=x^2

OpenStudy (anonymous):

I don't exactly understand the question, but I will clarify a few properties, which may help. y + 3 = 1/3^2 +1 This evaluates to: y = 1/3^2 + 1 - 2 y = 1/3^2 - 1 y = 1/9 - 1 Now replace y with 1/9 -1 So y = x^2 becomes 1/9 - 1 = x^2 x = + or - sqrt(1/9 - 1) Hope this offers some assistance.